Find X so that for any 3 × 3 real matrix A you get AX = XA = A [Hint : what property is being exhibited by real number p so that for any real w we get wp = pw = w then interpret for matrices.]
1
Expert's answer
2021-10-06T17:33:09-0400

Require to find "X" so that for any "3\times 3" real matrix "A" such that "AX=XA=A"

Recollect the following property in real numbers:


If "a" is any real number, then "a\cdot 1=a=1\cdot a" and 1 is called the multiplicative identity


Using the above, we have for any real number "w",

"wp=pw=w" then "p" is called the multiplicative identity and "p=1"


Now let us interpret the same for matrices


For any real matrix "A",

"AX=XA=A" then the matrix "X" is called the multiplicative identity and "X" is identity matrix or unit matrix.


Since "A" is "3\times 3" real matrix, so "X" is also "3\times 3" matrix

Therefore, "X=\begin{bmatrix}
1 & 0 & 0\\ 
 0& 1 & 0\\ 
0 & 0 & 1
\end{bmatrix}"